2. If you use a 1 µl inoculating loop to streak a plate from a urine specimen and count 140 colonies on the plate after incubating overnight, how much bacteria was in the original sample in CFU/ml (please write your answer in scientific notation, 1.1 x 106 for example?
Answer #1

1 ml = 1000 ul

1 ul contains 140 colonies

So colonies in 1 ml = 140 * 1000 = 14 * 104